OpenStudy (anonymous):

Choose the point-slope form of the equation below that represents the line that passes through the point (6, -3) and has a slope of 1/2 A. y - 6 = 1/2 (x + 3) B. y = 1/2x - 6 C. y + 3 = -1/5 (x - 2) D. y - 2 = -1/5 (x + 3)

OpenStudy (anonymous):

All I need is the answer, I'll give a medal! Thanks guys.

pooja195 (pooja195):

No.

OpenStudy (anonymous):

Thats your choice no medal though

pooja195 (pooja195):

I already helped. You are on a LEARNING site where we give HELP not answers.

OpenStudy (anonymous):

Please, I came on here to find help with some answers, don't worry this isn't homework or a test.

pooja195 (pooja195):

\(\color{blue}{\text{Originally Posted by}}\) @Abbyster Please, I came on here to find help with some answers, don't worry this isn't homework or a test. \(\color{blue}{\text{End of Quote}}\) Keyword: HELP you came for help

OpenStudy (anonymous):

Help finding ANSWERS

OpenStudy (anonymous):

y\[(y-y1)=m(x-x1)\]

OpenStudy (anonymous):

y+3=1/2 (x-6)

OpenStudy (anonymous):

y = 1/2x - 6

OpenStudy (anonymous):

B